public class Test {

private static int i =0;

private static int j =0;

public static void main(String[] args)

{

int i = 2;

int k = 3;

{

int j =3;

System.out.println("i + j is : " + i + j);

}

k = i + j;

System.out.println("K is " + k );

System.out.println("K is " + j);

}

}

why is the output i + j = 23

K =2

K =0

Please explain a step by step to get to the answer.
Answer #1

ANSWER:-

System.out.println("i + j is : " + i + j); // OUTPUT 23 because in this statement did not use ( ) so did not add i and j and print seprate value of i and like i=2 and y=3 means print 23

k = i + j;

System.out.println("K is " + k ); // OUTPUT 2 because i intialize in main function so scope of i work in full main function but j intialize in local scope in bracket so access globally means j=0

then k=2+0 print 2

System.out.println("K is " + j); // OUTPUT 0

j intialize in local scope in bracket so out of the bracket work globally which is initialize with 0 so access globally means j=0 so print 0

// if any doubt please comment and vote

then k=2+0 print 2